A solution is prepared by dissolving 19.2 g of glucose in 205 g of water. The final
volume of the solution is 214 ml. Find the concentration of the solution in units of
molality.
1
Expert's answer
2021-09-17T02:08:22-0400

250 ml × 0.4 = 100ml

= 100/2 = 50ml

Molar mass of sodium hydrogen carbonate = 84.007

Molarity = 111/84.007 = 1.32 mol

1.32/20

= 0.066mol/L